solve the following

(x+2)/(4) + (3x+1)/(3)
1
Expert's answer
2012-11-29T08:04:53-0500
(x+2)/(4) + (3x+1)/(3)=(3(x+2) +4(3x+1))/12=(3x+6+12x+4)/12=(15x+10)/12
If (15x+10)/12=0, then 15x+10=0
x=-2/3